OpenStudy (buggiethebug):

After t minutes of growth, a certain culture has a mass in grams of M(t)=2^t a) what is the average rate of growth during the time interval 3 ≤t≤3.00 Is the answer 5.547...?

OpenStudy (anonymous):

i think there may be a typo here, since \(3=3.00\)

OpenStudy (anonymous):

it cant really say \(3\leq t\leq 3.00\) right? that doesn't make sense

OpenStudy (anonymous):

unless it is asking for the instantaneous rate of change, rather than the average rate of change, in which case the answer would be \(8\ln(2)\)

OpenStudy (buggiethebug):

i meant 3.001

OpenStudy (buggiethebug):

from 3 ≤t≤3.001 sorry about the error

OpenStudy (anonymous):

Hint: \[\frac{d}{dt} a^t = (\ln a) \cdot a^t\]

OpenStudy (buggiethebug):

cant use derivatives... i just want to know if my answer is correct.